40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10796     Entry time: Sat Dec 13 14:26:36 2014
Author: ericq 
Type: Update 
Category: LSC 
Subject: Mismatched gains on ETMY Transmon QPD 

Yesterday, we were seeing anomalously high low frequency RIN in the y-arm (rms of 4% or so). I swung by the lab briefly to check this out. Turns out, despite TRY of 1.0, there was reasonable misalignment. ASS with the excitation lowered by a factor of two, and overall gain at 0.5 or so aligned things to TRY=1.2, and the RIN is back down to ~0.5% I reset the Thorlabs FM to make the power = 1.0

I then went to center the transmitted beam on the transmon QPD. Looking at the quadrant counts as I moved the beam around, things looked odd, and I poked around a little... 

I strongly suspect that we have significantly mismatched gains for the different quadrants on the ETMY QPD. 

Reasoning: With the y-arm POY locked, I used a lens to focus down the TRY beam, to illuminate the quadrants individually. Quadrants 2 and 3 would go up to 3 counts, while 1 and 4 would go up to 0.3 and 0.6, respectively. (These counts are in some arbitrary units that were set by setting the sum to 1.0 when pitch and yaw claimed to be centered, but mismatched gains makes that meaningless.)

I haven't looked more deeply into where the mismatch is occurring. The four individual whitening gain sliders did affect the signals, so the sliders don't seem sticky, however I didn't check the actual change in gains. Will the latest round of whitening board modifications help this?

Hopefully, once this is resolved, the DC transmission signals will be much more reliable when locking...

ELOG V3.1.3-